Computer Science Division

READER APPLICATIONS FOR PAID READER POSITIONS

(To work for CS300 Units, contact the instructor)

NAME:

(print)  Last  First  Middle Initial

ADDRESS: PHONE: NETWORK ADDRESS: BIRTHDATE: SOC. SEC. #: GPA: NUMBER OF UNITS ENROLLED THIS TERM: COURSE(S) YOU WOULD LIKE TO BE A READER FOR: WHEN DID YOU TAKE THE COURSES? GRADE(S) RECEIVED STATUS: GRADUATE UNDERGRADUATE PREVIOUS EMPLOYMENT: READER/TUTOR EXPERIENCE (including CS300)

(continue on back of sheet if needed)

COURSE:  INSTRUCTOR:  DATE:
COURSE:  INSTRUCTOR:  DATE:

CURRENT OR PREVIOUS CAMPUS EMPLOYMENT DEPARTMENT: POSITION: DEPARTMENT: POSITION: LANGUAGES AND COMPUTER SYSTEMS YOU’VE USED: On the back of this sheet, list any other Computer Science Courses you’ve taken. Please state the semester/year and the grades you’ve received. After completing this application, please return it to 571 Evans. Student’s signature/Date:

- - - - - - - - - - - - - - - - - - - - - - - - - - - - - - - - - - - - - - - - - - - - - - - - - - - - - - - - - - - - - - - - - - - - - - - - - - - - - - - -

This portion to be completed by instructor: The formula to determine reader hours/week is: enrollment x unit value of course (this may change – check with Gene Lawler), divided by 30. The hours allotted for the semester will change to reflect the true enrollment. Use ACE enrollment figures only as an estimate.

I have selected  to read for course  ˙
Course enrollment is  Ṫhere is/are  paid readers for this course.

Instructor’s signature/Date: